Please help me with this system of equations
2x-3y-z=4
4x+y+z=15
x-y-z=-2

Answer by stanbon(75887)

Please help me with this system of equations:
--------------------
2x-3y-z=4
4x+y+z=15
x-y-z=-2
=================
Rearrange:
x - y - z = -2
2x-3y - z = 4
4x+ y + z = 15
-----------------
Multiply 1st equation by 2 and subtract from 2nd equation
Multiply 1st equation by 4 and subtract from 3rd equation
----------------------
x - y - z = -2
0 - y + z = 8
0 + 5y +5z= 23
------------------------
Multiply 2nd equation by 5 and add to 3rd Equation
x - y - z = -2
0 - y + z = 8
0 + 0 + 10z = 63
-----
z = 6.3
----
Substitute and solve for "y":
-y + 6.3 = 8
y = -1.7
-----
Substitute and solve for "x":
x -(-1.7) - (6.3) = -2
x + 1.7 = 4.3
x = 2.6
=============================
Cheers,
Stan H.